# Oryx: a Performant and Scalable Algorithm for Many-Agent Coordination in Offline MARL

### v1

@@ 5 Related Work
Offline MARL. Early works such as Jiang, Lu (2021) and Yang et al. (2021) introduced methods to mitigate extrapolation errors through constrained Q-value estimation within the training distribution. Pan et al. (2022) combined first-order gradient methods with zeroth-order optimisation to guide policies toward high-value actions, with further refinements provided by Shao et al. (2023), which applied conservative regularisation using per-agent counterfactual reasoning, and Wang et al. (2023), employing a global-to-local value decomposition approach. Distributional constraints have been another promising direction, notably in works like Matsunaga et al. (2023) and Bui et al. (2025). Numerous additional works have also tackled extrapolation error, coordination issues, and offline stability (Zhang et al., 2023a; Wu et al., 2023; Eldeeb et al., 2024; Liu et al., 2024b; Barde et al., 2024; Liu et al., 2025; Zhou et al., 2025). Our approach builds directly upon these insights, specifically extending the early ICQ framework (Yang et al., 2021) to enhance learning stability from offline trajectories.
MARL as a Sequence Modeling Problem. Capturing long-term dependencies and joint-agent behaviour is critical, particularly in partially observable scenarios and tasks with extended horizons. Prior works such as Meng et al. (2022) and Tseng et al. (2022) addressed these challenges through Transformer-based architectures, effectively modeling trajectory data in offline MARL settings. Attention-based and diffusion-based strategies have also been explored (Zhu et al., 2025; Qiao et al., 2025; Li et al., 2025; Fu et al., 2025). Online MARL research further contributed with influential works like (Wen et al., 2022), which introduced transformer-based coordination mechanisms, alongside newer architectures aimed at better scalability with number of agents (Daniel et al., 2025; Mahjoub et al., 2025).
Moreover, theoretical advancements have enhanced understanding of the guarantees and limitations inherent to offline MARL (Cui, Du, 2022a, b; Zhong et al., 2022; Zhang et al., 2023b; Xiong et al., 2023; Wu et al., 2023). Complementary studies investigated opponent modeling in offline contexts (Jing et al., 2024), and explorations into offline-to-online transitions (Zhong et al., 2024a; Formanek et al., 2023b) reveal potential pathways to bridge static offline training with dynamic online adaptation.
@@ 6 Conclusion
In this work, we introduced Oryx, a novel algorithm explicitly designed to address the critical challenges of coordination amongst many agents in offline MARL. Our primary contribution is the derivation of a sequential policy updating scheme that leverages implicit constraint Q-learning (Yang et al., 2021) and the advantage decomposition theorem (Kuba et al., 2021). By integrating this sequential ICQ-based updating mechanism with a modified version of the Sable network (Mahjoub et al., 2025), Oryx effectively mitigates two fundamental problems in offline MARL: extrapolation error and miscoordination.
Our extensive empirical evaluation across diverse benchmarks—including SMAC, RWARE, Multi-Agent MuJoCo, and Connector—demonstrates that Oryx consistently achieves state-of-the-art results. Notably, Oryx excels in scenarios characterised by high agent densities and complex coordination requirements, distinguishing it from prior approaches for offline multi-agent learning. To help spur the field in this direction, we make all our datasets on Connector, with up to 50 agents, openly accessible to the community, a valuable contribution in and of itself.
Overall, Oryx represents a significant advancement in offline MARL and could provide useful insights for deploying coordinated multi-agent systems in real-world applications where environmental interaction is constrained or prohibitively expensive.
@@
While Oryx performs robustly across diverse research benchmarks with great potential for application, its evaluation remains far from the true complexity of offline MARL in large-scale real-world industrial settings. Future work may explore the extension of Oryx to hybrid offline-online settings and investigate its applicability and effectiveness across broader and more varied domains, including real-world applications.
@@ A.1 Proof of Theorem 1
For an auto-regressive model, the multi-agent joint-policy under ICQ regularisation can be optimised for $j=1,...,n$ as follows:
@@
$\mathcal{J}_{\boldsymbol{\pi}}$ $=\mathbb{E}_{\boldsymbol{\tau},\,\mathbf{a}\sim B}\Big{[}-\frac{1}{Z(\boldsymbol{\tau})}\log\bigl{(}\pi(\mathbf{a}\mid\boldsymbol{\tau})\bigr{)}\exp\!\Bigl{(}\tfrac{A(\boldsymbol{\tau},\mathbf{a})}{\alpha}\Bigr{)}\Big{]}$
$=\mathbb{E}_{\boldsymbol{\tau},\,\mathbf{a}^{i_{1:n}}\sim B}\Big{[}-\frac{1}{Z(\boldsymbol{\tau})}\Bigl{(}\sum_{j=1}^{n}\log\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}})\Bigr{)}\exp\!\Bigl{(}\tfrac{\sum_{l=1}^{n}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\Bigr{)}\Big{]}$
$=\sum_{j=1}^{n}\mathbb{E}_{\boldsymbol{\tau},\,\mathbf{a}^{i_{1:n}}\sim B}\Big{[}-\frac{1}{Z(\boldsymbol{\tau})}\log\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}})\exp\!\Bigl{(}\tfrac{\sum_{l=1}^{n}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\Bigr{)}\Big{]}$
$=\sum_{j=1}^{n}\mathbb{E}_{\boldsymbol{\tau},\,\mathbf{a}^{i_{1:j}},\,\mathbf{a}^{i_{j+1:n}}\sim B}\Bigl{[}-\frac{1}{Z(\boldsymbol{\tau})}\log\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}})\exp\!\Bigl{(}\tfrac{\sum_{l=1}^{j}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\Bigr{)}\Bigr{.}$
.$\cdot\exp\!\Bigl{(}\tfrac{\sum_{k=j+1}^{n}A^{i_{k}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:k-1}},a^{i_{k}})}{\alpha}\Bigr{)}\Bigr{]}$
$=\sum_{j=1}^{n}\mathbb{E}_{\boldsymbol{\tau},\,\mathbf{a}^{i_{1:j}}\sim B}\Big{[}-\frac{1}{Z^{i_{1:j}}(\boldsymbol{\tau})}\log\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}})\exp\!\Bigl{(}\tfrac{\sum_{l=1}^{j}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\Bigr{)}\Big{]}$
$\quad\cdot\mathbb{E}_{\boldsymbol{\tau},\,\mathbf{a}^{i_{1:n}}\sim B}\Big{[}\frac{1}{Z^{i_{j+1:n}}(\boldsymbol{\tau})}\exp\!\Bigl{(}\tfrac{\sum_{k=j+1}^{n}A^{i_{k}}(\boldsymbol{\tau},\mathbf{a}^{i_{j+1:k-1}},a^{i_{k}})}{\alpha}\Bigr{)}\Big{]}$
$=\sum_{j=1}^{n}\mathbb{E}_{\boldsymbol{\tau},\,\mathbf{a}^{i_{1:j}}\sim B}\Big{[}-\frac{1}{Z^{i_{1:j}}(\boldsymbol{\tau})}\log\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}})\exp\!\Bigl{(}\tfrac{A^{i_{1:j}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:j}})}{\alpha}\Bigr{)}\Big{]}$
@@ B.3.1 Implementation Details
Besides the baselines and Oryx, we test the effect of isolating Oryx’s key components. The first ablation concerns disabling auto-regressive actions, in the original implementation, agent $i+1$ receives the actions of agents $i,i-1,...,1$ as an input, within the same timestep, to support sequential coordination. To disable this mechanism, we instead feed a constant placeholder value of -1 as the previous action input to all agents during both training and evaluation. The second ablation targets removing temporal dependency, to achieve this, we reduce the sequence length from the default value of 20 (used in Oryx) to 2, such that only timesteps $t$ and $t+1$ are provided during training. This limited context prevents the model from capturing or leveraging long-term temporal dependencies, effectively disabling its ability to retain memory across steps. The third ablation is about removing the ICQ loss component, specifically, we eliminate the advantage estimation and policy regression terms from the loss function, retaining only standard Q-learning.
@@ C.1 Environment Details
The Connector environment (Bonnet et al., 2024) is a cooperative multi-agent grid world where each agent is randomly assigned a start and end position and must construct a path to connect the two. As agents move, they leave behind impassable trails, introducing the need for coordination to avoid blocking others. The action space is discrete with five options: up, down, left, right, and no-op. Agents observe a local $d\times d$ view centered on their position, including visible trails, their own coordinates, and all target locations. The reward function assigns $+1$ when an agent successfully connects to its target and $-0.03$ at every other timestep, with no further reward once connected.
@@ D.1 About SMAX
SMAX (Rutherford et al., 2024) is a JAX-based reimplementation of SMAC (Samvelyan et al., 2019a) and SMAC(v2) (Ellis et al., 2023), designed for efficient experimentation without the need for the StarCraft II engine. In this environment, agents form teams of heterogeneous units and cooperate to win battles in a real-time strategy setting. Each agent observes local information—such as positions, health, unit types, and recent actions of nearby allies and enemies—and selects from a discrete action space including movement and attack commands. Unlike SMAC, SMAX balances reward signals between tactical engagements (damage dealt) and final success (winning the episode).

### v2

@@ 5 Related Work
Offline MARL. Early works such as Jiang and Lu (2021) and Yang et al. (2021) introduced methods to mitigate extrapolation errors through constrained Q-value estimation within the training distribution. Pan et al. (2022) combined first-order gradient methods with zeroth-order optimisation to guide policies toward high-value actions, with further refinements provided by Shao et al. (2023), which applied conservative regularisation using per-agent counterfactual reasoning, and Wang et al. (2023), employing a global-to-local value decomposition approach. Distributional constraints have been another promising direction, notably in works like Matsunaga et al. (2023) and Bui et al. (2025). Numerous additional works have also tackled extrapolation error, coordination issues, and offline stability (Zhang et al., 2023a; Wu et al., 2023; Eldeeb et al., 2024; Liu et al., 2024b; Barde et al., 2024; Liu et al., 2025; Zhou et al., 2025). Our approach builds directly upon these insights, specifically extending the early ICQ framework (Yang et al., 2021) to enhance learning stability from offline trajectories.
Moreover, theoretical advancements have enhanced understanding of the guarantees and limitations inherent to offline MARL (Cui and Du, 2022b, a; Zhong et al., 2022; Zhang et al., 2023b; Xiong et al., 2023; Wu et al., 2023). Complementary studies investigated opponent modeling in offline contexts (Jing et al., 2024), and explorations into offline-to-online transitions (Zhong et al., 2024a; Formanek et al., 2023b) reveal potential pathways to bridge static offline training with dynamic online adaptation.
MARL as a Sequence Modeling Problem. Capturing long-term dependencies and joint-agent behaviour is critical, particularly in partially observable scenarios and tasks with extended horizons. Prior works such as Meng et al. (2022) and Tseng et al. (2022) addressed these challenges through Transformer-based architectures, effectively modeling trajectory data in offline MARL settings. Attention-based and diffusion-based strategies have also been explored (Zhu et al., 2024; Qiao et al., 2025; Li et al., 2025; Fu et al., 2025). Online MARL research further contributed with influential works like by Wen et al. (2022), which introduced a transformer-based autoregressive action selection mechanism, alongside newer architectures aimed at better scalability with number of agents (Daniel et al., 2025; Mahjoub et al., 2025).
@@ 6 Conclusion
In this work, we introduced Oryx, a novel algorithm explicitly designed to address the critical challenges of coordination amongst many agents in offline MARL. We derived a sequential policy updating scheme that leverages implicit constraint Q-learning (Yang et al., 2021) and the advantage decomposition theorem (Kuba et al., 2021). By integrating this sequential ICQ-based updating scheme with a modified version of the Sable network (Mahjoub et al., 2025), Oryx effectively mitigates two fundamental problems in offline MARL: extrapolation error and miscoordination. Our extensive empirical evaluation across diverse benchmarks—including SMAC, RWARE, Multi-Agent MuJoCo, and Connector—demonstrates that Oryx consistently achieves state-of-the-art results. Notably, Oryx excels in scenarios characterised by high agent densities and complex coordination requirements, distinguishing it from prior approaches for offline multi-agent learning. To help accelerate research in this direction, we make all our datasets on Connector, with up to 50 agents, openly accessible to the community.
@@
While Oryx demonstrates robust performance across diverse research benchmarks, its evaluation is naturally limited compared to the true complexity of large-scale, real-world industrial settings. Consequently, important future work includes extending Oryx to hybrid offline-online settings and evaluating its effectiveness in broader, more varied domains, particularly real-world applications. Furthermore, our work introduces autoregressive policies to offline MARL, demonstrating the significant promise of utlising such policies in other architectures and setups. Future research could extend existing and novel offline MARL methods to utilise autoregressive policies (Fu et al., 2022) and sequence models. Integrating diverse offline learning techniques into this paradigm may lead to the discovery of even more effective sequence models for multi-agent coordination.
Acknowledgments
We would like to thank Louise Beyers for her valuable contributions at the outset of this project, which were crucial for its inception. We thank our MLOps team for developing our model training and experiment orchestration platform AIchor. We thank the Python and JAX communities for developing tools that made this research possible. We thank the anonymous reviewers for their constructive feedback and valuable suggestions. Finally, we thank Google’s TPU Research Cloud (TRC) for supporting our research with Cloud TPUs.
@@ A.1 Proof of Theorem 1
For an auto-regressive model, the multi-agent joint-policy under ICQ regularisation can be optimised for j=1,,nj=1,...,n as follows:
@@
$\mathcal{J}_{\boldsymbol{\pi}}$ $=\mathbb{E}_{\boldsymbol{\tau},\mathbf{a}\sim B}\left[-\frac{1}{Z(\boldsymbol{\tau})}\log(\pi(\mathbf{a}|\boldsymbol{\tau}))\exp\left(\frac{A(\boldsymbol{\tau},\mathbf{a})}{\alpha}\right)\right]$
$=\mathbb{E}_{\boldsymbol{\tau},\mathbf{a}^{i_{1:n}}\sim B}\left[-\frac{1}{Z(\boldsymbol{\tau})}\left(\sum^{n}_{j=1}\log(\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}}))\right)\exp\left(\frac{\sum_{l=1}^{n}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\right)\right]$
$=\sum^{n}_{j=1}\mathbb{E}_{\boldsymbol{\tau},\mathbf{a}^{i_{1:n}}\sim B}\left[-\frac{1}{Z(\boldsymbol{\tau})}\left(\log(\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}}))\right)\exp\left(\frac{\sum_{l=1}^{n}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\right)\right]$
$=\sum^{n}_{j=1}\mathbb{E}_{\boldsymbol{\tau},\mathbf{a}^{i_{1:j}},\mathbf{a}^{i_{j+1:n}}\sim B}\left[-\frac{1}{Z(\boldsymbol{\tau})}\log(\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}}))\exp\left(\frac{\sum_{l=1}^{j}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\right)\right$.
[…]
$=\sum^{n}_{j=1}\mathbb{E}_{\boldsymbol{\tau},\mathbf{a}^{i_{1:j}}\sim B}\left[-\frac{1}{Z^{i_{1:j}}(\boldsymbol{\tau})}\log(\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}}))\exp\left(\frac{\sum_{l=1}^{j}A^{i_{l}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:l-1}},a^{i_{l}})}{\alpha}\right)\right]$
$\hskip 50.00008pt\cdot\mathbb{E}_{\boldsymbol{\tau},\mathbf{a}^{i_{1:n}}\sim B}\left[\frac{1}{Z^{i_{j+1:n}}(\boldsymbol{\tau})}\exp\left(\frac{\sum_{k=j+1}^{n}A^{i_{k}}(\boldsymbol{\tau},\mathbf{a}^{i_{j+1:k-1}},a^{i_{k}})}{\alpha}\right)\right]$
$=\sum^{n}_{j=1}\mathbb{E}_{\boldsymbol{\tau},\mathbf{a}^{i_{1:j}}\sim B}\left[-\frac{1}{Z^{i_{1:j}}(\boldsymbol{\tau})}\log(\pi^{i_{j}}(a^{i_{j}}\mid\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}}))\exp\left(\frac{A^{i_{1:j}}(\boldsymbol{\tau},\mathbf{a}^{i_{1:j}})}{\alpha}\right)\right]$
@@ B.3.1 Implementation Details
Besides the baselines and Oryx, we test the effect of isolating Oryx’s key components. The first ablation concerns disabling auto-regressive actions, in the original implementation, agent $i+1$ receives the actions of agents i,i1,,1i,i-1,...,1 as an input, within the same timestep, to support sequential coordination. To disable this mechanism, we instead feed a constant placeholder value of -1 as the previous action input to all agents during both training and evaluation. The second ablation targets removing temporal dependency, to achieve this, we reduce the sequence length from the default value of 20 (used in Oryx) to 2, such that only timesteps $t$ and $t+1$ are provided during training. This limited context prevents the model from capturing or leveraging long-term temporal dependencies, effectively disabling its ability to retain memory across steps. The third ablation is about removing the ICQ loss component, specifically, we eliminate the advantage estimation and policy regression terms from the loss function, retaining only standard Q-learning.
@@ C.1 Environment Details
The Connector environment (Bonnet et al., 2024) is a cooperative multi-agent grid world where each agent is randomly assigned a start and end position and must construct a path to connect the two. As agents move, they leave behind impassable trails, introducing the need for coordination to avoid blocking others. The action space is discrete with five options: up, down, left, right, and no-op. Agents observe a local $d\times d$ view centered on their position, including visible trails, their own coordinates, and all target locations. The reward function assigns +1+1 when an agent successfully connects to its target and 0.03-0.03 at every other timestep, with no further reward once connected.
@@ D.1 About SMAX
SMAX (Rutherford et al., 2024) is a JAX-based reimplementation of SMAC (Samvelyan et al., 2019) and SMAC(v2) (Ellis et al., 2023), designed for efficient experimentation without the need for the StarCraft II engine. In this environment, agents form teams of heterogeneous units and cooperate to win battles in a real-time strategy setting. Each agent observes local information—such as positions, health, unit types, and recent actions of nearby allies and enemies—and selects from a discrete action space including movement and attack commands. Unlike SMAC, SMAX balances reward signals between tactical engagements (damage dealt) and final success (winning the episode).